OpenStudy (anonymous):

will reward and medal. just need two equations let x= first number let x=2nd number equations 55.)find two consecutive odd integers whose sum is 628?? 56.)find three consecutive integers whose sum is -1,230 57.)find three consecutive integers whose sum is -198 58.)find three consecutive odd integers whose sum is 339 59.)find three consecutive even integers whose sum is 342?

jimthompson5910 (jim_thompson5910):

55.)find two consecutive odd integers whose sum is 628 let x = first number so x+2 is the next consecutive odd integer adding them and setting the sum equal to 628 gives us x+x+2 = 628 2x+2 = 628 2x = 628-2 2x = 626 x = 626/2 x = 313 x+2 = 313+2 =315 Therefore, the two integers are 313 and 315

OpenStudy (anonymous):

thanks @jimthompson

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (anonymous):

@jimthompson would u like to have medal or no??? your choice

jimthompson5910 (jim_thompson5910):

either way is fine

OpenStudy (anonymous):

help w/next one just need two equations u use to solve???? i can rest

jimthompson5910 (jim_thompson5910):

find three consecutive integers whose sum is -1,230 let x = first integer, so x+1 is the next consecutive integer and x+2 is the next one add them up, set them equal to -1230 to get x+x+1+x+2 = -1230

OpenStudy (anonymous):

3x+5=1230

OpenStudy (anonymous):

right? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

close

jimthompson5910 (jim_thompson5910):

x+x+1+x+2 = -1230 3x+3 = -1230

jimthompson5910 (jim_thompson5910):

keep going

OpenStudy (anonymous):

x=409 for first x what do i do for 2nd x and 3rd x?

jimthompson5910 (jim_thompson5910):

3x+3 = -1230 3x = -1230-3 3x = -1233 x = -1233/3 x = -411

jimthompson5910 (jim_thompson5910):

x = -411 x+1 = -411+1 = -410 x+2 = -411+2 = -409 so the three integers are: -411, -410, -409

OpenStudy (anonymous):

thanks. help w/ next one???

OpenStudy (anonymous):

I may be able to help u with the next one, once you share the problem with me.

jimthompson5910 (jim_thompson5910):

I'll set it up, but let you finish

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

okay jim then you can take over. I won't bump in.

jimthompson5910 (jim_thompson5910):

57.)find three consecutive integers whose sum is -198 x = first number x+1 = next consecutive number x+2 = next consecutive number add up the 3 numbers and set it equal to -198 x+x+1+x+2 = -198 solve for x and tejeshwar95, you can help as much as you want

OpenStudy (anonymous):

well, okay... but it seems to me that you got this. but will ask away whenever u need help

OpenStudy (anonymous):

how do u do the just x one/????

OpenStudy (anonymous):

1st step on equation??????

OpenStudy (anonymous):

Just add the three numbers that is x + x +1 + x + 2 = -198

OpenStudy (anonymous):

When you ask for the first step, that is where the question ends.

OpenStudy (anonymous):

practically speaking.

OpenStudy (anonymous):

if i could id give u both a best response u guys have been sooo much help

OpenStudy (anonymous):

who would rather have a best response more??? if u care

OpenStudy (anonymous):

Do u want another question like this for practice sake?

OpenStudy (anonymous):

58 and 59 idk how to set up three equations that's all i need?

OpenStudy (anonymous):

Give it to the other guy, he is the main guy, I merely butt in so.... I don't really deserve it.

OpenStudy (anonymous):

couldone of u hep with 58 and 59??

OpenStudy (anonymous):

Okay, this is really easy. Assume a number x

OpenStudy (anonymous):

now count the distance between two odd numbers lets say the odd numbers are 1 and 3

OpenStudy (anonymous):

now try to form the equation. Its okay to do it wrong, but its important to do it.

OpenStudy (anonymous):

please @tejeshaw95 look at how @jimthompson explained it u don't make any sense to me?????

OpenStudy (anonymous):

if u just gave me equations w/out explaining it. that's NOT cheating

OpenStudy (anonymous):

okay. Lemme explain it in a bit more detail. I was actually hoping you would take the hint and run with it. But, its okay. well, lets assume that the odd numbers we are talking about are 1,3,5. Now, these are what are called consecutive odd numbers. now see the difference between two consecutive odd numbers. 3-1 = 2 5-3 = 2 Now we all know that 1+3 + 5 is not = to 339. So, now, lets assume the first number to be x. So, the number following it would be x + 2 since that is the difference or gap between two consecutive odd numbers. And the odd number after the 2nd odd number which was x+2 is x+2+2 which is x+4. So, now we add em we get x+ (x+2)+(x+4) now, the sum of these three numbers is 339. So, now u solve the equation.

OpenStudy (anonymous):

Consecutive means one after the other. so three consecutive numbers can be 1,2,3. Two consecutive even numbers can be 0,2,4 and two consecutive odd numbers can be 1,3,5.

OpenStudy (anonymous):

3x+1+5=339 ??????

OpenStudy (anonymous):

2nd equation above could it be written that way??

OpenStudy (anonymous):

you are technically wrong but, right. actually it should've been 3x+2+4

OpenStudy (anonymous):

Lemme give you something easier to help u understand the concept. Find three consecutive numbers whose sum is 6.

OpenStudy (anonymous):

I got 111 please help w/1st and third equation. 2+2+2=6 Are u allowed to just give me 1st and third equation????

OpenStudy (anonymous):

If I give you the equation, the question reaches its end. The trick lies in forming the equation. No, consecutive doesn't mean the same numbers. It has to be three different numbers.

OpenStudy (anonymous):

2+3+1

OpenStudy (anonymous):

yes. that is correct.

OpenStudy (anonymous):

how does that help w/58???? or 59

OpenStudy (anonymous):

because it is the basic concept that allows you to solve the question. You are too focused on finding the answer to the question to such an extent that you are missing out the underlying concept.

OpenStudy (anonymous):

I am giving you the concept so u can solve millions of such problems without any help, and trust me that is a great feeling.

OpenStudy (anonymous):

I know the problem and how many answers u need to find. and I know how to solve equations. I just don't know what the equations are. I understand what ur saying. its just like when jimjohnson helped iwas learning and starting to understand he gave me equations and step by step and I was starting to get it.. now u wont imstarting to not comprehend as well

OpenStudy (anonymous):

I have 5 more questions after I get 58 and 59 understood

OpenStudy (anonymous):

please reply????? @teleshaw95

OpenStudy (anonymous):

@tejeshwar95 plz reply?

OpenStudy (anonymous):

Oh, I see what you want. 58.)find three consecutive odd integers whose sum is 339 let x = first number so x+2 is the next consecutive odd integer and x+4 is the odd integer after that adding them and setting the sum equal to 628 gives us x + (x+2) + (x+4) = 339 x + x + x + 2 + 4 = 339 3x = 339-2-4 3x = 333 x = 111 x = 111 x +2 = 111+2 =113 x + 4 = 111 + 4 =115

OpenStudy (anonymous):

oh btw u got x right. Now you just gotta add the 2 for the second number and 4 for the third number. Sorry, my bad, I hadn't calculated the answer then.

OpenStudy (anonymous):

that explaining was perfect all I need is equation 1st line tho per equation.. 59??? do same way pleaseeeeee???????

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

asap pleaseeee

OpenStudy (anonymous):

59.)find three consecutive even integers whose sum is 342? let x = first number so x+2 is the next consecutive even integer and x+4 is the even integer after that adding them and setting the sum equal to 342 gives us x + (x+2) + (x+4) = 342 x + x + x + 2 + 4 = 342 3x = 342-2-4 3x = 336 x = 112 x +2 = 112 +2 =114 x + 4 = 112 + 4 =116

OpenStudy (anonymous):

I got 112 and 116 what did x=?

OpenStudy (anonymous):

x = 112, x+2 = 114, x +4 = 116